5) A 6.0 kg object is experiencing 2 forces with identical magnitude of 10.0 N. What is the maximum and minimum magnitude of the acceleration that the object can experience based on the possible directions that the forces may act upon the object?
Answer #1

F1 = F2 = 10 N

mass = m = 6 kg

maximum force=F = F1 + F2 = 10 + 10 = 20 N

maximum acceleration = a = F / m = 20 / 6 = 3.33 m/sec^2

minimum force = F1 - F2 = 10 - 10 = 0

minimum acceleration = 0/6 = 0 m/sec^2